OpenStudy (anonymous):

Prove that the left side equals the right side. (Sinx/1-cosx) + (sinx/1+cosx) = 2 csc x

ganeshie8 (ganeshie8):

take LHS, pull out Sinx, take LCM simplify

OpenStudy (anonymous):

What is LHS?

ganeshie8 (ganeshie8):

Left Hand Side

ganeshie8 (ganeshie8):

(Sinx/1-cosx) + (sinx/1+cosx)

ganeshie8 (ganeshie8):

pull out Sinx

OpenStudy (anonymous):

Ok so I've got sinx(1/cos+1)(1/cos-1). What do you mean take LCM

ganeshie8 (ganeshie8):

\(\large \sin x (\frac{1}{1-\cos x} + \frac{1}{1 + \cos x})\)

OpenStudy (anonymous):

OOOHHHH. sec=1/cos Thanks I see it now.

ganeshie8 (ganeshie8):

\(\large \sin x (\frac{1}{1-\cos x} + \frac{1}{1 + \cos x})\) \(\large \sin x (\frac{1+\cos x + 1 - \cos x}{1-\cos^2 x})\)

ganeshie8 (ganeshie8):

easy for u :)